ROK-C issued a 90-day commercial papers (CPs) with the face value of $2,050,000. If these CPs are sold at $1,900,000. The 90-day interest payment is $____; the 90-day interest rate is _____%; and its EAR on these commercial papers is ______%.
Answer #1

90 day interest payment = Face value - Price

= 2050000-1900000

=$150,000

90 day interest rate = 150000/2050000 = 7.32%

EAR = (1+APR)^n -1

= (1+ 7.32%)^4 -1

=32.65%